Solve using any method.
3x – 11y = 9
-9x + 33y = 18

Answer by Edwin McCravy(20056)

 3x – 11y = 9
-9x + 33y = 18

By elimination:

Multiply the first equation through by 3,
and add the 2nd equation to it term by terms:

 9x – 33y = 27
-9x + 33y = 18
--------------
 0x +  0y = 45

There is no value of x or y that can make this equation true,
as the left side will always be 0, and the right side 45.
So there can be no solution.  This is what is called an
"inconsistent system". If you try to solve it graphically,
you get these two parallel lines:



which never intersect and therefore can give no solution.

Edwin